OpenStudy (anonymous):

the area of a square region is 18 m square. the length of its diagonal is? Please also explain how you solve it.

OpenStudy (anonymous):

use the Pythagorean theorem.

OpenStudy (anonymous):

or the triangle property: \(1-1-\sqrt{2}\)\[\]

OpenStudy (anonymous):

finding the square's side from its area should be simple.

OpenStudy (anonymous):

I don't know about the Triangle property. How would i use Pythagorian theorem in it? No side is given?

OpenStudy (anonymous):

\[A_{square}=side^2\]

OpenStudy (anonymous):

Diagonal of a square is always and always is \( \sqrt{2} \times \text{ side of the square} \)

OpenStudy (turingtest):

if a side is x then\[A=x^2\] right? and from the pythagorean theorem we get \[h^2=2x^2\]so we can find h between the two equations